Session: Nov 1998 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Maria J. Rioja at Wed Nov 25 01:18:18 MET 1998) Minor failures
Failure in head position at the beginning of first pass on 3rd tape. The error was fixed few minutes afterwards, by ca. 19:05.

Jodrell (Alastair Gunn at Wed Nov 25 07:50:05 MET 1998) Success
No known problems.

Cambridge (Alastair Gunn at Wed Nov 25 08:20:02 MET 1998) Minor failures
Possible problem with head stack positions after 22:38 UT. Requires investigation after schedule end. Data may be recoverable.

Torun (K.Borkowski at Wed Nov 25 10:00:46 MET 1998) Minor failures
Due to technical problems recordings are useful only from the scan at 8:09 UT. The scan at 22:09 partly (till 22:19) recorded off-source (since antenna had to rotate through 300 degrees in the azimuth).

Medicina (orfei at Wed Nov 25 10:37:30 MET 1998) Minor failures
Addition to comment of the operator. As third tape a THIN tape(!) was used instead of a thick one. Anyway it will be sent to Bonn correlator.

Noto (C.Trigilio at Thu Nov 26 10:11:42 MET 1998) Minor failures
Lost one hour from 18:50 to 19:40 due to power failure at the VLBA terminal.

Onsala (A. Polatidis at Thu Nov 26 10:12:53 MET 1998) Success
No problems with the observations

Effelsberg ( at Thu Nov 26 13:48:45 MET 1998) Success
seems OK

Westerbork (Tony Foley at Tue Dec 1 18:07:28 MET 1998) Minor failures
Formatter clock jumped from +0.35usec to -0.45usec at about 12UT, for no obvious reason. The maser clock did not jump. The array needed rephasing twice.

Calibration (Jean\-Francois Desmurs at Fri Dec 11 16:30:21 MET 1998) Success
(Ef overflow on VC09). The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: ec007.antab. Read the comments in the file to get further information.